OpenStudy (anonymous):

Hah I Have A Stupid Question.. But Honestly I Have No Idea What It Is Haha... :) What is x^2 + x ? My Mind Shut Off. Would It be x^3 ? But wouldn't That be If I multiplied the x ? Thank You!!

OpenStudy (anonymous):

\(x^2\times x=x^3\) just like \(100\times 10=1000\) but you cannot combine the terms you have

OpenStudy (anonymous):

you can factor it if you like and rewrite it as \(x(x+1)\) but there is not particular reason to do that

OpenStudy (anonymous):

Yes!! The factoring part.. I knew I was missing something :) thank you @satellite73 !!

OpenStudy (anonymous):

yw